OpenStudy (anonymous):

Find all polar coordinates of point P where P = (4.-pi/3.)

OpenStudy (jdoe0001):

hint: all co-terminal points to \(\left( 4,-\frac{\pi }{3} \right)\) 

OpenStudy (anonymous):

is that right @jdoe0001

OpenStudy (jdoe0001):

hmmm well how would you get a 2nd coterminal point how much would you go around to get a 2nd terminal point?

OpenStudy (anonymous):

i really dunno

OpenStudy (anonymous):

can you help me? @jdoe0001

OpenStudy (jdoe0001):

well.... do you know what a coterminal point is?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok.. so... how many radians would you need to go around to get the 1st coterminal point? 

OpenStudy (jdoe0001):

in order to get a 2nd co-terminal point, you'd need to do a "full turn", or \(\large 2\pi\) the "4" is just the length of the radius, so that stays steady, only thing is that, it could also be negative so.... now, notice your choices "n" stands for an integer, thus \(\large \begin{array}{llll} coterminal&degrees \\\hline\\ 1&-\frac{\pi }{3}+2\pi\\ 2&-\frac{\pi }{3}+4\pi\\ 3&-\frac{\pi }{3}+6\pi\\ 4&-\frac{\pi }{3}+8\pi\\ ...&... \end{array}\)

OpenStudy (anonymous):

ohhhh, okay i think i get it! @jdoe0001

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

thank you! @jdoe0001